Bethesda Mining Company reports the following balance sheet information for 2018 and 2019.
BETHESDA MINING COMPANY
Balance Sheets as of December 31, 2018 and 2019
2018 2019 2018 2019
Assets Liabilities and Owners’ Equity
  Current assets   Current liabilities
    Cash $ 55,526 $ 70,205      Accounts payable $ 188,922 $ 196,611
    Accounts receivable 63,281 83,639      Notes payable 84,020 135,588
    Inventory 121,382 186,805
            Total $ 272,942 $ 332,199
      Total $ 240,189 $ 340,649
  Long-term debt $ 235,000 $ 171,750
  Owners’ equity
     Common stock and paid-in surplus $ 220,000 $ 220,000
     Accumulated retained earnings 170,594 206,478
  Fixed assets   
    Net plant and equipment $ 658,347 $ 589,778            Total $ 390,594 $ 426,478
  Total assets $ 898,536 $ 930,427   Total liabilities and owners’ equity $ 898,536 $ 930,427

Calculate the following financial ratios for each year:  a.Current ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)b.Quick ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)c.Cash ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)d.Debt-equity ratio and equity multiplier. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)e.Total debt ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

a. 2018 current ratio times
2019 current ratio times
b. 2018 Quick ratio times
2019 Quick ratio times
c. 2018 Cash ratio times
2019 Cash ratio times
d. 2018 Debt-equity ratio times
2018 Equity multiplier times
2019 Debt-equity ratio times
2019 Equity multiplier times
e 2018 Total debt ratio times
2019 Total debt ratio times
Answer #1

Part a)

The current ratio is calculated as below:

Current Ratio = Total Current Assets/Total Current Liabilities

Using the values provided in the question, we get,

2018 Current Ratio = 240,189/272,942 = .88 times

2019 Current Ratio = 340,649/332,199 = 1.03 times

______

Part b)

The value of quick ratio for each year is arrived as follows:

Quick Ratio = (Total Current Assets - Inventory)/Total Current Liabilities

Using the values provided in the question, we get,

2018 Quick Ratio = (240,189-121,382)/272,942 = .44 times

2019 Quick Ratio = (340,649-186,805)/332,199 = .46 times

______

Part c)

The value of cash ratio for each year is determined as below:

Cash Ratio = Cash/Current Liabilities

Using the values provided in the question, we get,

2018 Cash Ratio = 55,526/272,942 = .20 times

2019 Cash Ratio = 70,205/332,199 = .21 times

______

Part d)

The value of debt-equity ratio and equity multiplier for each year is calculated as follows:

Debt-Equity Ratio = Total Debt/Total Equity

Equity Multiplier = 1+ Debt-Equity Ratio

Using the values provided in the question, we get,

2018 Debt-Equity Ratio = (272,942+235,000)/390,594 = 1.30 times

2018 Equity Multiplier = 1.30+1 = 2.30 times

2019 Debt-Equity Ratio = (332,199+171,750)/426,478 = 1.18 times

2019 Equity Multiplier = 1.18+1 = 2.18 times

______

Part e)

The value of total debt ratio for each year is calculated as below:

Total Debt Ratio = Total Debt/Total Assets

Using the values provided in the question, we get,

2018 Total Debt Ratio = (272,942+235,000)/898,536 = .57 times

2019 Total Debt Ratio = (332,199+171,750)/930,427 = .54 times

______

Tabular Representation:

a. 2018 Current ratio 0.88 times
2019 Current ratio 1.03 times
b. 2018 Quick ratio 0.44 times
2019 Quick ratio 0.46 times
c. 2018 Cash ratio 0.20 times
2019 Cash ratio 0.21 times
d. 2018 Debt-equity ratio 1.30 times
2018 Equity multiplier 2.30 times
2019 Debt-equity ratio 1.18 times
2019 Equity multiplier 2.18 times
e 2018 Total Debt ratio 0.57 times
2019 Total Debt ratio 0.54 times